# Achieving designed texture and flows in bulk active nematics using optimal control theory

### v1

@@ I INTRODUCTION
The precise control of fluids is essential for biological processes including motility and material transport across cell, tissue and organismal scales of organization [1, 2, 3, 4, 5, 6, 7]. These flows are generated and modulated by ‘active’ protein structures that convert chemical energy into mechanical forces and motion. While these protein sources are generated at the molecular (nm) scale, they self-organize to generate and control fluid flows on much larger (micron) scales. For example, cytoskeletal networks composed of filaments and motor proteins induce cytoplasmic flows or streams within cells that drive processes ranging from chloroplast transport in algae to the recycling of motility proteins during eukaryotic chemotaxis [8, 9, 10, 11, 12, 13]. Notably, control and self-regulation are essential to target the cytoskeleton to a particular functional dynamical state, selected from the rich manifold of available states [14, 15, 16, 17, 18, 19]. Understanding the mechanisms of biological fluid control will enable developing synthetic systems that use similar principles of self-organization to enable macroscale control of flows, thus creating advanced functional, responsive, and adaptive materials [20, 21].
@@
Existing uncontrolled active matter systems composed of purified filaments and motor proteins exhibit nonequilibrium, self-organized structures including microtubule asters [22, 23, 24, 25, 26], contractile microtubule networks [22, 23], and can generate spontaneous fluid flows [27, 28, 29, 30]. However, these flows are typically chaotic and disorganized, and thus not capable of producing work or useful functions. One approach to organize active matter flows is to use confining boundaries [31, 32, 30, 33, 34, 35, 36, 37, 38, 39]. However, a more versatile approach, typically employed by biological systems, is to implement feedforward and/or feedback control. Experiments have recently taken a key step toward this approach by demonstrating that light can be used to control the spatiotemporal patterns or assembly of active colloids [40, 41] and microswimmers [42, 4, 43, 44, 45, 46, 47]. Similarly, researchers have recently used optogenetic kinesin motor proteins to develop light-activated microtubule-based active nematic systems, enabling controlling the average flow speed [48, 49], steering defects [50, 51], and driving the assembly and motions of asters [52]. However, despite these remarkable successes, in general it remains unclear how to choose a spatiotemporal sequence of light that will drive the system into a predetermined state with useful functions. This inverse problem is particularly challenging because active matter undergoes an energy cascade from forces generated at the particle scale to mesoscale active stresses that ultimately drive the emergent dynamics. Thus, more robust control protocols are required.
@@
Recently, theoretical works have taken two approaches toward controlling active matter into functionalized states. In the first, investigators subject systems to predetermined activity patterns, and observe the resulting system behaviors to identify possible target state dynamics [53, 50, 54, 55, 56, 57, 58]. In the second, researchers choose a predetermined target state, and use techniques such as optimal control theory [59, 60, 61, 62, 53, 63, 64, 65, 66], optimal transport [67], or reinforcement learning [68, 69, 70] applied to a continuum hydrodynamic model for the active matter system to identify spatiotemporal activity patterns that drive the system toward the target. More broadly, optimal control theory has been used to find control laws for dynamical systems modeled by partial differential equations (PDEs) and systems such as Newtonian fluids [71, 72], traffic flows [73], designing optimal interplanetary trajectories [74], or to identify optimal trajectories for individual active colloids [75, 76].
@@
More closely related to the present article, recent works used optimal control theory to respectively switch the chirality of the steady state in circularly confined active nematics [62] or alter the direction of or switch between stable attractors (stationary asters and propagating strips) in a polar active fluid [61]. While those results demonstrate the capability to transition between symmetric states that are attractors in the absence of control, here we show that optimal control protocols can be applied much more broadly, driving the system into states that cannot be obtained or stabilized without control, and do not correspond to exact coherent structures [77, 78]. Other works have applied protocols obtained through trial-and-error [79], optimal control solutions [53, 62] or symmetry considerations to perform tasks such as steering individual defects [80]. Here, we show that optimal control theory can perform significantly more complex tasks, such as eliminating an ensemble of defects and then maintaining a defect-free region indefinitely. The challenging nature of this task is evident from the complexity of the spatiotemporal activity pattern of this control solution, in particular the activity gradients required to prevent defects from entering the boundaries of the square region of interest that we chose. Recently, Serra et al. [81] showed that one can use optimal control to regulate material transport (e.g. creating material traps) by generating short-time attractors. While that approach has advantages in some situations, our approach is straightforward to implement for virtually any type of emergent behavior.
@@ II MODEL
$\partial_{t}\mathbf{Q}+\nabla\cdot(\mathbf{u}\mathbf{Q})+[\mathbf{Q},\mathbf{\Omega}]-\lambda\mathbf{E}-\mathbf{H}$
@@
where $\lambda=1$ is the flow-aligning parameter, $\eta=1$ is the coefficient of viscosity, $\xi=0.01$ is the strength of substrate friction, the mechanical pressure $P$ is such that flows are incompressible: $\nabla\cdot\mathbf{u}=0$, and $\Omega_{ij}=\frac{1}{2}\left(\partial_{i}u_{j}-\partial_{j}u_{i}\right)$ and $E_{ij}=\frac{1}{2}\left(\partial_{i}u_{j}+\partial_{j}u_{i}\right)$ are, respectively, the anti-symmetric and symmetric parts of the flow field gradient. We added the small substrate friction to aid convergence of the control solutions; because the associated frictional screening length is much larger than the active length scale for all relevant parameters in this work, it does not qualitatively affect our results. The parameter $\alpha$ is the activity ($\alpha>0:$ extensile, $\alpha<0:$ contractile). The H term describes the relaxational dynamics of the nematic tensor to the minimum of a free energy through a molecular field defined as
@@ III OPTIMAL CONTROL
Finally, we solve iteratively for the optimal control variable $\alpha(\bm{x},t)$ by coupling the Stokes equations with the associated adjoint system. The conditions for optimality are $(\delta\mathcal{L}/\delta\psi),(\delta\mathcal{L}/\delta\nu)$, $(\delta\mathcal{L}/\delta\phi),\quad(\delta\mathcal{L}/\delta\mathbf{Q}),\quad(\delta\mathcal{L}/\delta\mathbf{u}),\quad(\delta\mathcal{L}/\delta P),\quad(\delta\mathcal{L}/\delta\alpha)=0$. The first three conditions return the original nemato-hydrodynamics equations governing $\mathbf{Q}$, $\mathbf{u}$, and $P$. The following two conditions yield the dynamical equations for the adjoint variables $\bm{\psi}$, $\bm{\nu}$, and $\phi$
@@
We use the direct-adjoint-looping (DAL) method [82] to minimize the cost function under the constraint that the dynamics satisfies Eqs. (1) and (2), to yield the optimal schedule of activity in space and time $\alpha(\bm{x},t)$. Specifically, we construct an initial condition corresponding to the stable steady state at the parameter set of interest, and specify a time duration $t_{\text{F}}$ over which the control protocol will be employed, and an initial trial control protocol. We then perform a series of DAL iterations; in each iteration the system dynamics are integrated from the initial condition $\mathbf{Q}(\bm{x},0),\mathbf{u}(\bm{x},0),\alpha(\bm{x},0)$ for time $t_{\text{F}}$ under the current control protocol, and the cost function (5) is computed from the resulting trajectory. The adjoint equations are integrated backwards in time, which propagates the residuals. After each backward run, the control protocol is updated via gradient descent, $\alpha^{i+i}=\alpha^{i}-\delta\mathcal{L}/\delta\alpha$ using Eq. (6). We employ Armijo backtracking [83] to compute the step-size of gradient descent and to ensure convergence of the DAL algorithm. We have implemented this calculation in the open source Python finite element method solver FEniCS.
Figure 3: Aligned director field in a square domain. The target state has a uniform alignment of the director field at an orientation of $\pi/4$ within the square highlighted region. The director field is unconstrained outside of this region, and there are no constraints on the velocity field anywhere. (a,b) Snapshots from a representative trajectory under the optimal control protocol with (a) the nematic order parameter $s$ (color map) and director field orientation (lines) and (b) the control solution activity field $\alpha$ (color map), with brighter/darker colors representing higher/lower activity. Topological defects are indicated by color, with $+1/2$ in green and $-1/2$ in cyan. The parameters used in this optimization are $\{\alpha_{0},J_{1},J_{2},J_{3},\zeta,W_{\text{u}},W_{\text{Q}},t_{\text{F}}\}=\{0.0,1.0,1.0,0.1,0.1,0.0,100,100.0\}$. (c,d,e) Statistics of the system behavior inside and outside the square region: (c) mean activity, (d) defect density, and (e) total elastic energy.
@@ IV RESULTS
Fig. 2 shows the behavior of the bulk system in the absence of control. We see that for values of activity below a threshold ($\alpha^{*}\lesssim 0.17$) the system undergoes bend fluctuations but does not form defects on simulated timescales. Above this threshold, we observe spontaneous defect nucleation, with a density $\propto 1/l_{\text{a}}^{2}$, with the active length scale $l_{\text{a}}\cong\sqrt{K/\alpha}$ as expected from theory [84]. Similarly, the mean vortex size decreases with $\alpha$. Importantly, these states are the only behaviors that we observe above and below the threshold activity value, showing that control is essential to achieve other emerging patterns from this system.
@@ IV.1 Nematic alignment of a subdomain within a turbulent bulk active nematic
With our first target state, we demonstrate controllability of nematic tensor field, $\mathbf{Q}$ (Fig. 3). Starting from an initial condition corresponding to a turbulent bulk active nematic, we aim to ‘flatten’ the nematic within a subdomain (highlighted square region in Fig. 3). Specifically, the target state has the nematic uniformly aligned along the $\pi/4$ axis within the subdomain, and no constraints on the director field elsewhere. There are no constraints on the velocity profile $W_{\text{u}}=0$ for this target. To obtain the initial condition, we start with a nearly uniformly aligned nematic (with 1 % random noise) and zero velocity, and evolve the system with Eqs. (1) and (2) under uniform activity ($\alpha=2.5$) until it reaches steady state. This results in an initial configuration with 15 defects in the subdomain. We then compute the optimal control solution for $t_{\text{F}}=100$.
@@
For further insight into the forces that drive flattening of the nematic, we present characteristics of the system behaviors inside and outside of the subdomain as a function of time during the course of the control window (Fig. 3c,d,e). Fig. 3c shows the mean activity as a characteristic of the strength of the control input. We see that the mean activity is relatively constant over time, and significantly smaller inside the subdomain compared to the exterior at earlier times. Over time mean activity of the exterior decreases as well. We then show the defect density in Fig.3c, average total elastic energy ($1/2(\mathbf{\nabla Q})^{2}$) in Fig. 3d, and deviation from the target state ($\Delta\mathbf{Q}:\Delta\mathbf{Q}$) in SI 8a to characterize the rate at which the solution converges. Interestingly, we see that the elastic energy rapidly decreases both inside and outside the subdomain during the early stage of the trajectory, but reaches a significantly smaller value within the aligned region. The fact that the elastic energy decreases outside of the subdomain, even though this region plays no role in the target state configuration or cost function, can be attributed to the fact that defects are slowed and annihilated as they approach the subdomain, thus decreasing the total defect density and elastic energy of the sample. This effect can be attributed to the finite size of the sample; in a larger sample there would be a finite correlation length $\propto\sqrt{K/\langle\alpha\rangle}$ beyond which influence from the subdomain would decay. Finally, Fig. 3e shows that the mean flow rate $u$ is roughly constant over time once the domain has flattened, but that flow rates are higher outside than inside the domain. This difference is primarily caused by the higher mean activity outside, but also by the the defects in the exterior, which generate active stress imbalances that can increase the flow rate.

### v2

@@ I Introduction:
The precise control of fluids is essential for biological processes including motility and material transport across cell, tissue and organismal scales of organization Fletcher and Mullins (2010); Hawkins et al. (2009); Berg (2004); Liebchen and Löwen (2018); Pierce-Shimomura et al. (2008); Videler (1993); Behkam and Sitti (2007); Angelini et al. (2011); Duclos et al. (2018). These flows are generated and modulated by ‘active’ protein structures that convert chemical energy into mechanical forces and motion. While these protein sources are generated at the molecular (nm) scale, they self-organize to generate and control fluid flows on much larger (micron) scales. For example, cytoskeletal networks composed of filaments and motor proteins induce cytoplasmic flows or streams within cells that drive processes ranging from chloroplast transport in algae to the recycling of motility proteins during eukaryotic chemotaxis Allen and Allen (1978); Keren et al. (2009); Monteith et al. (2016); Keller and Segel (1971). Notably, control and self-regulation are essential to target the cytoskeleton to a particular functional dynamical state, selected from the rich manifold of available states Blanch-Mercader et al. (2018); Poujade et al. (2007); Serra-Picamal et al. (2012). Understanding the mechanisms of biological fluid control will enable developing synthetic systems that use similar principles of self-organization to enable macroscale control of flows, thus creating advanced functional, responsive, and adaptive materials Shao and Fu (2014); Qu et al. (2021); Wu et al. (2017).
@@
Existing uncontrolled active matter systems composed of purified filaments and motor proteins exhibit nonequilibrium, self-organized structures including microtubule asters Foster et al. (2015); Najma et al. (2024); Berezney et al. (2022); Hentrich and Surrey (2010); Ndlec et al. (1997), contractile microtubule networks Foster et al. (2015); Najma et al. (2024), and spontaneous fluid flows Sanchez et al. (2012); DeCamp et al. (2015); Ellis et al. (2018); Lushi, Wioland, and Goldstein (2014). However, these flows are typically chaotic and disorganized, and thus not capable of producing work or useful functions. One approach to organize active matter flows is to use confining boundaries Varghese et al. (2020); Wu et al. (2017); Wioland et al. (2013); Lushi, Wioland, and Goldstein (2014); Hardoüin et al. (2019); Opathalage et al. (2019); Uplap, Hagan, and Baskaran (2023); Fily, Baskaran, and Hagan (2014, 2015); Fazli and Naji (2021); Iyer, Gompper, and Fedosov (2023). However, a more versatile approach, typically employed by biological systems, is to implement feedforward and/or feedback control. Experiments have recently taken a key step toward this approach by demonstrating that light can be used to control the spatiotemporal patterns or assembly of active colloids Aubret et al. (2018); Palacci et al. (2013) and microswimmers Volpe et al. (2011); Liebchen and Löwen (2018); Huang, Omori, and Ishikawa (2020); Winkler and Gompper (2020); Villa and Pumera (2019); Qi et al. (2022); Katuri et al. (2017); Kurzthaler et al. (2024). Similarly, researchers have recently used optogenetic motor proteins to develop light-activated active nematic systems, enabling controlling the average flow speed Lemma et al. (2023); Zarei et al. (2023); Najma et al. (2022); Zhang et al. (2021), steering defects Zhang et al. (2021); Repula et al. (2024); Zhang, Snezhko, and Sokolov (2022), and driving the assembly and motions of asters Ross et al. (2019). However, despite these remarkable successes, in general it remains unclear how to choose a spatiotemporal sequence of light that will drive the system into a predetermined state with useful functions. This inverse problem is particularly challenging because active matter undergoes an energy cascade from forces generated at the particle scale to mesoscale active stresses that ultimately drive the emergent dynamics. Thus, more robust control protocols are required.
@@
Recently, theoretical works have taken two approaches toward controlling active matter into functionalized states. In the first, investigators subject systems to predetermined activity patterns, and observe the resulting system behaviors to identify possible target state dynamics Shankar et al. (2022); Zhang et al. (2021); Nasiri, Löwen, and Liebchen (2023); Knežević, Welker, and Stark (2022); Rønning, Marchetti, and Angheluta (2023); Desgranges et al. (2023); Zhang, Snezhko, and Sokolov (2022). In the second, researchers choose a predetermined target state, and use techniques such as optimal control theory Brunton and Kutz (2022); Dullerud and Paganini (2013); Ghosh et al. (2024); Norton et al. (2020); Shankar et al. (2022); Sinigaglia, Braghin, and Serra (2023); Mowlavi and Nabi (2023); Garcia-Millan et al. (2024); Grover, Bakshi, and Theodorou (2018), optimal transport Villani et al. (2009), or reinforcement learning Floyd, Dinner, and Vaikuntanathan (2024a); Chennakesavalu et al. (2024); Falk et al. (2021) applied to a continuum hydrodynamic model for the active matter system to identify spatiotemporal activity patterns that drive the system toward the target. More broadly, optimal control theory has been used to find control laws for dynamical systems modeled by partial differential equations (PDEs). These efforts include modeling systems ranging from Newtonian fluids Nabi, Grover, and Caulfield (2017); Passaggia and Ehrenstein (2010) to traffic flows Gugat et al. (2005), designing optimal interplanetary trajectories Liu and Barabási (2016), and identifying optimal trajectories for individual active colloids Schneider and Stark (2019); Liebchen and Löwen (2019).
@@
More closely related to the present article, recent works used optimal control theory to respectively switch the chirality of the steady state in circularly confined active nematics Norton et al. (2020) or alter the direction of or switch between stable attractors (stationary asters and propagating strips) in a polar active fluid Ghosh et al. (2024); Shankar et al. (2022). While those results demonstrate the capability to transition between symmetric states that are attractors in the absence of control, here we show that optimal control protocols can be applied much more broadly, driving the system into states that cannot be obtained or stabilized without control, and do not correspond to exact coherent structures Wagner et al. (2022, 2023). Other works have applied protocols obtained through trial-and-error Floyd, Dinner, and Vaikuntanathan (2024b), optimal control solutions Shankar et al. (2022); Norton et al. (2020) or symmetry considerations to perform tasks such as steering individual defects Shankar et al. (2024). Here, we show that optimal control theory can perform significantly more complex tasks, such as eliminating an ensemble of defects and then maintaining a defect-free region indefinitely. The challenging nature of this task is evident from the complexity of the spatiotemporal activity pattern of this control solution, in particular the activity gradients required to prevent defects from entering the boundaries of the square region of interest that we chose. Recently, Serra et al. Sinigaglia, Braghin, and Serra (2024) showed that one can use optimal control to regulate material transport (e.g. creating material traps) by generating short-time attractors. While that approach is very powerful for specific situations that are globally well-controlled, our approach is straightforward to implement and works well for emergent behavior using local control that is robust to global noise.
@@ II Model
$\partial_{t}\mathbf{Q}+\nabla\cdot(\mathbf{u}\mathbf{Q})+[\mathbf{Q},\bm{\Omega}]-\lambda\mathbf{E}-\mathbf{H}$
@@
where $\lambda=1$ is the flow-aligning parameter, $\eta=1$ is the coefficient of viscosity, $\xi=0.01$ is the strength of substrate friction, the mechanical pressure $P$ is such that flows are incompressible: $\nabla\cdot\mathbf{u}=0$, and $\Omega_{ij}=\frac{1}{2}\left(\partial_{i}u_{j}-\partial_{j}u_{i}\right)$ and $E_{ij}=\frac{1}{2}\left(\partial_{i}u_{j}+\partial_{j}u_{i}\right)$ are, respectively, the anti-symmetric and symmetric parts of the flow field gradient. We added the small substrate friction to aid convergence of the control solutions; because the associated frictional screening length is much larger than the active length scale for all relevant parameters in this work, it does not qualitatively affect our results Doostmohammadi et al. (2016); Alert, Martínez-Calvo, and Datta (2022); Chaithanya et al. (2024); Schimming, Reichhardt, and Reichhardt (2024a); Nejad, Doostmohammadi, and Yeomans (2021). The parameter $\alpha$ is the activity ($\alpha>0:$ extensile, $\alpha<0:$ contractile). The H term describes the relaxational dynamics of the nematic tensor to the minimum of a free energy through a molecular field defined as
@@
Figure 3: Aligned director field in a square domain. The target state has a uniform alignment of the director field at an orientation of $\pi/4$ within the square highlighted region. The director field is unconstrained outside of this region, and there are no constraints on the velocity field anywhere. (a,b) Snapshots from a representative trajectory under the optimal control protocol with (a) the nematic order parameter $s$ (color map) and director field orientation (lines) and (b) the control solution activity field $\alpha$ (color map), with brighter/darker colors representing higher/lower activity. Topological defects are indicated by color, with $+1/2$ in green and $-1/2$ in cyan. The parameters used in this optimization are $\{\alpha_{0},J_{1},J_{2},J_{3},\zeta,W_{\text{u}},W_{\text{Q}},t_{\text{F}}\}=\{0.0,1.0,1.0,0.1,0.1,0.0,100,100.0\}$. (c,d,e) Statistics of the system behavior inside and outside the square region: (c) mean activity $\langle\alpha\rangle$, (d) defect density, and (e) total elastic energy $\frac{1}{2}K(\mathbf{\nabla Q})^{2}$ normalized by spatial extent. A video of this trajectory is in supplemental Movie S1 SIr .
@@ III Optimal Control
Finally, we solve iteratively for the optimal control variable $\alpha(\bm{x},t)$ by coupling the Stokes equations with the associated adjoint system. The conditions for optimality are $(\delta\mathcal{L}/\delta\psi),(\delta\mathcal{L}/\delta\nu)$, $(\delta\mathcal{L}/\delta\phi),\quad(\delta\mathcal{L}/\delta\mathbf{Q}),\quad(\delta\mathcal{L}/\delta\mathbf{u}),\quad(\delta\mathcal{L}/\delta P),\\ \quad(\delta\mathcal{L}/\delta\alpha)=0$. The first three conditions return the original nemato-hydrodynamics equations governing $\mathbf{Q}$, $\mathbf{u}$, and $P$. The following three conditions yield the dynamical equations for the adjoint variables $\bm{\psi}$, $\bm{\nu}$, and $\phi$
@@
We use the direct-adjoint-looping (DAL) method Kerswell, Pringle, and Willis (2014) to minimize the cost function under the constraint that the dynamics satisfies Eqs. (1) and (2), to yield the optimal schedule of activity in space and time $\alpha(\bm{x},t)$. Specifically, we construct an initial condition corresponding to the stable steady state at the parameter set of interest, and specify a time duration $t_{\text{F}}$ over which the control protocol will be employed, and an initial trial control protocol. We then perform a series of DAL iterations; in each iteration the system dynamics are integrated from the initial condition $\mathbf{Q}(\bm{x},0),\mathbf{u}(\bm{x},0),\alpha(\bm{x},0)$ for time $t_{\text{F}}$ under the current control protocol, and the cost function (5) is computed from the resulting trajectory. The adjoint equations are integrated backwards in time, which propagates the residuals. After each backward run, the control protocol is updated via gradient descent, $\alpha^{i+i}=\alpha^{i}-\delta\mathcal{L}/\delta\alpha$ using Eq. (6). We employ Armijo backtracking Borzì and Schulz (2011) to compute the step-size of gradient descent and to ensure convergence of the DAL algorithm. We have implemented this calculation in the open source Python finite element method solver FEniCS Dupont et al. (2003).
@@ IV Results
Fig. 2 shows the behavior of the bulk system in the absence of control. We see that for values of activity below a threshold ($\alpha^{*}\lesssim 0.17$) the system undergoes small bend fluctuations but does not form defects on simulated timescales. This can be attributed to the long time scale required for defects to spontaneously form at low activity Giomi et al. (2011, 2012, 2013); Marchetti et al. (2013); Saintillan and Shelley (2013); Saintillan (2018); Theillard and Saintillan (2019). Above this threshold, we observe spontaneous defect nucleation, with a density $\propto 1/l_{\text{a}}^{2}$ and mean vortex size $\propto l_{\text{a}}$, with the active length scale $l_{\text{a}}\cong\sqrt{K/\alpha}$ as expected from theory Marchetti et al. (2013). Importantly, these states are the only behaviors that we observe above and below the threshold activity value, showing that control is essential to achieve other emerging patterns from this system.
@@
Figure 4: Coherent flow through a ‘channel’ without confining boundaries. The target state has uniform flow along the $\hat{x}$ direction with magnitude $U_{\text{T}}=0.5$ within the strip enclosed by green dashed lines. (a,e) The velocity field (arrows) and magnitude $u$ (color map) are shown at indicated time points for representative trajectories under the optimal control protocol, for the systems with y-dimension (a) $L_{y}=50$ and (b) $L_{y}=100$. (b,f) The activity field $\alpha$ (color map) and nematic director field (lines) are shown at the same time points for the two system sizes. The initial state has the nematic uniformly aligned along the $\pi/4$ axis in each case. (c,g) The flow speed $\langle\mathbf{u}_{\text{x}}\rangle_{x,t}$ and activity $\langle\alpha\rangle_{x,t}$ averaged over time and the channel axis ($\hat{x}$ direction) for the small and large system sizes respectively. (d, h) the flow speed $\langle\mathbf{u}_{\text{x}}\rangle_{x,y}$ averaged over the channel and bulk regions as a function of time for the two system sizes. The parameter values for this solution were $\{J_{1},J_{2},J_{3},\zeta,W_{\text{u}},W_{\text{Q}},\alpha_{0},t_{\text{F}}\}=\{1.0,1.0,0.1,0.1,80.0,0,0.1,200.0\}$. Videos corresponding to Videos of these trajectories are provided in supplemental movies S2 and S3 SIr .
@@ IV.1 Nematic alignment of a subdomain within a turbulent bulk active nematic
With our first target state, we demonstrate controllability of the nematic tensor field, $\mathbf{Q}$ (Fig. 3). Starting from an initial condition corresponding to a turbulent bulk active nematic, we aim to ‘flatten’ the nematic within a subdomain (highlighted square region in Fig. 3). Specifically, the target state has the nematic uniformly aligned along the $\pi/4$ axis within the subdomain, and no constraints on the director field elsewhere. There are no constraints on the velocity profile $W_{\text{u}}=0$ for this target. To obtain the initial condition, we start with a nearly uniformly aligned nematic (with 1 % random noise) and zero velocity, and evolve the system with Eqs. (1) and (2) under uniform activity ($\alpha=2.5$) until it reaches steady state. This results in an initial configuration with 15 defects in the subdomain. We then compute the optimal control solution for $t_{\text{F}}=100$.
@@
Figure 5: Couette flow in a channel without a moving boundary. (a) The velocity field (arrows) and magnitude $u$ (color map) are shown for three time points. (b) The control solution activity field $\alpha$ (color map) and nematic director field (lines) for the same time points.(c) Plot of $u_{x}$ (blue line, left y-axis) compared to the target profile (black dashed line, left y-axis), and the activity (red line, right y-axis) as a function of channel height at the final time point. The parameter values for this solution were $\{J_{1},J_{2},J_{3},\zeta,W_{\text{u}},W_{\text{Q}},t_{\text{F}}\}=\{8.0,1.0,0.1,0.1,15.0,0,15.0\}$. A video of this trajectory is in supplemental Movie S4SIr .
For further insight into the forces that drive flattening of the nematic, we present characteristics of the system behaviors inside and outside of the subdomain as a function of time during the course of the control window (Fig. 3c,d,e). Fig. 3c shows the mean activity as a characteristic of the strength of the control input. We see that the mean activity is relatively constant over time, and significantly smaller inside the subdomain compared to the exterior at earlier times. Over time mean activity of the exterior decreases as well. We then show the defect density in Fig.3d, average total elastic energy ($1/2K(\mathbf{\nabla Q})^{2}$) in Fig. 3e, and deviation from the target state ($\Delta\mathbf{Q}:\Delta\mathbf{Q}$) in SI 10a to characterize the rate at which the solution converges. Interestingly, we see that the elastic energy rapidly decreases both inside and outside the subdomain during the early stage of the trajectory, but reaches a significantly smaller value within the aligned region. The fact that the elastic energy decreases outside of the subdomain, even though this region plays no role in the target state configuration or cost function, can be attributed to the fact that defects are slowed and annihilated as they approach the subdomain, thus decreasing the total defect density and elastic energy of the sample. This effect can be attributed to the finite size of the sample; in a larger sample there would be a finite correlation length $\propto\sqrt{K/\langle\alpha\rangle}$ beyond which influence from the subdomain would decay.